OpenStudy (anonymous):

|3x-6| less than or equal too 21

OpenStudy (anonymous):

Less than what is 6x3?=18

OpenStudy (nbarrera):

It is less than 21

OpenStudy (anonymous):

I'll write "<=" to mean less than or equal to. |3x-6| <= 21 is the same as: -21 <= 3x - 6 <= 21, this is one of the rules of absolute value. Then you just have to solve for x, so add 6 to each side first: -21 + 6 <= 3x - 6 + 6 <= 32 + 6 -15 <= 3x <= 9 Divide by 3 to isolate the x, -5 <= x <= 9 So the solution to |3x-6| <= 21 is -5 <= x <= 9.

OpenStudy (anonymous):

Sorry about that typo in there --> the 32 should be 21: so when you're adding the 6: -21 + 6 <= 3x - 6 + 6 <= 21 + 6 -15 <= 3x <= 27 Then the rest is as I posted earlier.